OpenStudy (anonymous):

Directions: Type the correct answer in each box. Use numerals instead of words. If necessary, use / for the fraction bar(s). Expand the following expressions.

TheSmartOne (thesmartone):

Expression??

TheSmartOne (thesmartone):

Distibution aah

TheSmartOne (thesmartone):

So a(b-c)=ab-ac

TheSmartOne (thesmartone):

a here is -2 b is 3x c is 5

TheSmartOne (thesmartone):

So -2(3x-5)=-2*3x -( -2*5) =??

TheSmartOne (thesmartone):

* means multiplication let me write it in \(\ \LaTeX\)

TheSmartOne (thesmartone):

So what does this equal \(\Large (-2 \times 3x)- (-2 \times 5)\)